# Phase noise measurement of semiconductor optical amplifiers

### v1

@@ VI Second experiment
The second experiment aims at assessing the DUT white phase noise of $S_{\psi}(f)$ in different conditions of input power ($P_{i}=0.8\ldots 15\leavevmode\nobreak\ \upmu\mathrm{W}$) and pump current ($I_{P}=150\ldots 600\leavevmode\nobreak\ \mathrm{mA}$). We underline that in this Section we consider only the white phase noise $S_{\psi}$ and $S_{\varphi}$, thus ‘$(f)$’ is omitted.
Figure 4 shows the results for $I_{P}=350\leavevmode\nobreak\ \mathrm{mA}$. Notice that the horizontal axis is $P_{\text{RF}}$, instead of $P_{0}$. The reason is that $P_{\text{RF}}$ is measured directly, while $P_{0}$ can only be estimated using (7). In fact, a power meter measures the total power, which is $P_{1}+P_{2}+P_{\text{ASE}}$.
Figure 4: White phase noise of the DUT measured at $I_{P}=350\leavevmode\nobreak\ \mathrm{mA}$ pump power, for different values of the optical power $P_{i}$ at the DUT input. From the right hand to the left, the RF power is reduced by detuning the attenuator at the photodetector input. The background noise is shown for comparison.
@@
    Set the DUT input power $P_{i}$ to a convenient value. This changes $P_{0}$ without breaking the symmetry because the DUT is kept below the saturation point.
@@
The same measurement is repeated for 150 mA and 600 mA pump current. A synthesis of the results is shown in Fig. 5. The invalid points, $S_{\varphi}$ hidden under the background, are omitted. For this reason, at $I_{P}=150\leavevmode\nobreak\ \mathrm{mA}$ we can show only one point. Notice that the optical gain gets higher at higher $I_{P}$, thus the experimental points are limited to lower $P_{i}$.
Fig. 5 tells us that (i) no effect of the pump current $I_{P}$ is observed, and (ii) the DUT white noise is $S_{\varphi}\propto 1/P_{i}$, which is the phenomenological behavior of the shot noise. The latter point matches the obvious facts that optical signals do not suffer from thermal noise, and shot noise is inherent in the stimulated emission.
@@ VII Conclusions
We have discussed in detail a novel method for the phase-noise measurement of optical amplifiers using the delayed self-heterodyne intereferometric technique, and we have measured the phase noise of a semiconductor optical amplifier at 1.55 $\upmu\mathrm{m}$. Taking $-32\leavevmode\nobreak\ \mathrm{dBrad^{2}/Hz}$ at $f=1\leavevmode\nobreak\ \mathrm{Hz}$ as an upper bound of the flicker noise, the amplifier noise limits the fractional frequency stability to $5.2{\times}10^{-17}/\tau$ Allan deviation, as a function of the integration time $\tau$. This is lower than the instability of most Fabry-Pérot cavity-stabilized lasers [49, 50]. The flicker ($1/f$) phase noise could not be measured because it is of the same order or below the background. Operating the optical amplifier under different conditions of input power $P_{i}$ and pump current, we observed that the white phase noise is is independent of the pump current, and it is proportional to $1/P_{i}$, as expected for shot noise, and that it down to $-125$ $\mathrm{dBrad^{2}/Hz}$ at $f\geq 100$ $\mathrm{kHz}$.
Figure 5: Phase noise floor of the optical amplifier versus the SOA input optical power $P_{i}$, for several values (600, 350 and 150 mA) of the SOA driving current $I$. Only 1 point was taken for $I$ = 150 mA. In this case, the output RF signal was too low for lower values of $P_{i}$.

### v3

@@ VI Second experiment
The second experiment aims at assessing the DUT white phase noise of $S_{\psi}(f)$ in different conditions of input power ($P_{i}=0.8\ldots 15\penalty 10000\ \upmu\mathrm{W}$) and pump current ($I_{P}=150\ldots 600\penalty 10000\ \mathrm{mA}$). We underline that in this Section we consider only the white phase noise $S_{\psi}$ and $S_{\varphi}$, thus ‘$(f)$’ is omitted.
Figure 4 shows the results for $I_{P}=350\penalty 10000\ \mathrm{mA}$. Notice that the horizontal axis is $P_{\text{RF}}$, instead of $P_{0}$. The reason is that $P_{\text{RF}}$ is measured directly, while $P_{0}$ can only be estimated using (7). In fact, a power meter measures the total power, which is $P_{1}+P_{2}+P_{\text{ASE}}$.
Figure 4: White phase noise of the DUT measured at $I_{P}=350\penalty 10000\ \mathrm{mA}$ pump power, for different values of the optical power $P_{i}$ at the DUT input. From the right hand to the left, the RF power is reduced by detuning the attenuator at the photodetector input. The background noise is shown for comparison.
@@
    Set the DUT input power $P_{i}$ to a convenient value. Higher $P_{i}$ results in higher signal-to-noise ratio, but $P_{i}$ must be kept well below the saturation level of the DUT and of the photodetector. Keeping the amplifier in the linear region ensures that the symmetry condition is not affected by $P_{0}$.
@@
The same measurement is repeated for 150 mA and 600 mA pump current. A synthesis of the results is shown in Fig. 5. The invalid points, $S_{\varphi}$ hidden under the background, are omitted. For this reason, at $I_{P}=150\penalty 10000\ \mathrm{mA}$ we can show only one point. Notice that the optical gain gets higher at higher $I_{P}$, thus the experimental points are limited to lower $P_{i}$.
Figure 5 tells us that (i) no effect of the pump current $I_{P}$ is observed, and (ii) the DUT white noise is $S_{\varphi}\propto 1/P_{i}$, which is the phenomenological behavior of the shot noise. The latter point matches the obvious facts that optical signals do not suffer from thermal noise, and shot noise is inherent in the stimulated emission.
@@ VII Conclusions
We have discussed in detail a novel method for the phase-noise measurement of optical amplifiers using the delayed self-heterodyne intereferometric technique, and we have measured the phase noise of a semiconductor optical amplifier at 1.55 $\upmu\mathrm{m}$, operated below saturation. The method is suitable to laser amplifiers of other technologies, chiefly the EDFA, but the setup may need to be adapted (output power, gain, etc.).
Taking $32\penalty 10000\ \mathrm{dBrad^{2}/Hz}$ at $f=1\penalty 10000\ \mathrm{Hz}$ as an upper bound of the flicker noise, the amplifier noise limits the fractional frequency stability to $5.2{\times}10^{-17}/\tau$ Allan deviation, as a function of the integration time $\tau$. This is lower than the instability of most Fabry-Pérot cavity-stabilized lasers [50, 51]. The flicker ($1/f$) phase noise could not be measured because it is of the same order or below the background. Operating the optical amplifier under different conditions of input power $P_{i}$ and pump current, we observed that the white phase noise is (i) independent of the pump current, and (ii) proportional to $1/P_{i}$, as expected for shot noise, down to 126-126 $\mathrm{dBrad^{2}/Hz}$ at $f\geq 100$ $\mathrm{kHz}$.
Figure 5: Phase noise floor of the optical amplifier versus the SOA input optical power $P_{i}$, for several values (600, 350 and 150 mA) of the SOA driving current II. Only 1 point was taken for II = 150 mA. In this case, the output RF signal was too low for lower values of $P_{i}$.